# Multi-Robot Object SLAM using Distributed Variational Inference

### v2

@@ I INTRODUCTION
  • We demonstrate multi-robot object SLAM using stereo camera measurements for odometry and object detection on both real and simulated data to show that our method improves the overall accuracy of the trajectories and object maps of robot teams, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM [^1] Code is available at https://github.com/intrepidChw/distributed_msckf..
@@ II RELATED WORK
II-A Multi-robot bluefactor graph optimization
@@ II-B Multi-robot filtering
Compared to existing work, our contribution is to derive a fully distributed filtering approach for landmark-based SLAM from a constrained variational inference perspective. Our formulation makes a novel connection to distributed mirror descent and enables robots to achieve landmark consensus efficiently with one-hop communication only and without sharing private trajectory information.
@@ III PROBLEM STATEMENT
The robots communicate over a network represented as a connected undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $\mathcal{V}=\{1,\ldots,n\}$ corresponding to the robots and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ specifying robot pairs that can exchange information. For example, $(i,j)\in\mathcal{E}$ indicates that robot $i$ and $j$ can exchange information. Let $A\in\mathbb{R}^{n\times n}$ be a doubly stochastic weighted adjacency matrix of $\mathcal{G}$ such that $A_{ij}>0$ if $(i,j)\in\mathcal{E}$ and $A_{ij}=0$ otherwise. Also, let $\mathcal{N}_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}\cup\{i\}$ denote the set of one-hop neighbors of robot $i$. This is the set of all nodes $j\in\mathcal{V}$ that have a direct connection via an edge to node $i$ and includes node $i$ itself. We consider the following problem.
@@ IV-C Linear Gaussian case
As shown in [37, Chapter 3.3], considering only two consecutive time steps in the lifted form in (11) leads to a Kalman filter. Using the result in Proposition 3, we obtain a distributed Kalman filter that incorporates the consensus averaging step in (15). In [36, Appendix C], we also derive a version of the Kalman filter that allows correlation between the motion and measurement noise.
@@ V Distributed MSCKF
In this section, we use Proposition 3 to derive a distributed version of the MSCKF algorithm [8], summarized in Algorithm 1. Each step of the algorithms is described in the following subsections.
@@ V-B Consensus averaging
which is the same as (15) except that we only consider one time step $t$ here. Then, we need to reconstruct the new joint distribution $\mathcal{N}(\bar{\boldsymbol{\mu}}_{i,t-1},\bar{\Sigma}_{i,t-1})$. Since we store the historical camera poses as $SE(3)$ matrices, Proposition 2 can not be applied directly. Following [37, Chapter 7.3.1], we define a Gaussian distribution over a historical camera pose $\underline{T}_{i,k},\ k=t-c,\cdots,t-1$ by adding a perturbation $\epsilon_{i,k}$:
@@
where $(\cdot)^{\wedge}$ defined in [37, Chapter 7.1.2] converts from $\mathbb{R}^{6}$ to a $\mathbb{R}^{4\times 4}$ twist matrix. The estimated covariance $\Sigma_{i,t-1}$ already takes account of both the poses and landmarks. For consistency of notation with Proposition 2, we denote the mean of the pose perturbation as $\boldsymbol{\mu}^{\mathbf{x}}_{i,t-1}=\mathbf{0}_{6c}$. After averaging and reconstructing the new joint distribution, $\bar{\boldsymbol{\mu}}^{\mathbf{x}}_{i,t-1}$ may be non-zero, so we need to correct the camera poses as follows:
@@ V-C State propagation
where $Ad(\cdot)$ is the adjoint of an $SE(3)$ matrix [37, Chapter 7.1.4], $\mathbf{e}_{6n}\in\mathbb{R}^{6n+3m_{t}}$ is a vector with the $6n$-th element as $1$ and the rest as 00, and $W_{i}\in\mathbb{R}^{6\times 6}$ is the odometry measurement covariance.
@@ V-D State update
Finally, by concatenating the residuals and Jacobians of both geometric and object features, we get the overall residual $\mathbf{r}_{i,t}\!=\![\mathbf{r}^{g\top}_{i,t}\ \mathbf{r}^{o\top}_{i,t}]^{\top}$ and Jacobian $H_{i,t}\!=\![H^{g\top}_{i,t}\ H^{o\top}_{i,t}]^{\top}$. As shown in  [36, Appendix C], the Kalman gain is:
@@ A-C Derivation of the Kalman filter with noise correlation
Following the approach in Eq. 3.113 in [37, Chapter 3.3.2], we multiply both sides of (50) by the matrix:
@@
which aligns with [53, Table 5.1].

### v3

@@ Abstract
Index Terms:
Multi-Robot SLAM, Distributed Robot Systems, Probability and Statistical Methods.
@@ I INTRODUCTION
  • We demonstrate multi-robot object SLAM using stereo camera measurements for odometry and object detection on both real and simulated data to show that our method improves the overall accuracy of the trajectories and object maps of robot teams, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM [^1] Code is at https://github.com/ExistentialRobotics/distributed_msckf..
@@ II RELATED WORK
II-A Multi-robot factor graph optimization
@@ II-B Multi-robot filtering
Our contribution is to derive a fully distributed filter for object SLAM from a constrained variational inference perspective. Our formulation makes a novel connection to distributed mirror descent and enables robots to achieve landmark consensus efficiently with one-hop communication only and without sharing private trajectory information.
@@ III PROBLEM STATEMENT
The robots communicate over a network represented as a connected undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $\mathcal{V}=\{1,\ldots,n\}$ corresponding to the robots and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ specifying robot pairs that can exchange information, e.g. $(i,j)\in\mathcal{E}$ indicates that robot $i$ and $j$ can exchange information. Let $A\in\mathbb{R}^{n\times n}$ be a doubly stochastic weighted adjacency matrix of $\mathcal{G}$ such that $A_{ij}>0$ if $(i,j)\in\mathcal{E}$ and $A_{ij}=0$ otherwise. Also, let $\mathcal{N}_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}\cup\{i\}$ denote the set of one-hop neighbors of robot $i$.  and includes node $i$ itself. We consider the following problem.
@@ IV-C Linear Gaussian case
As shown in [36, Ch. 3.3], considering only two consecutive time steps in the lifted form in (11) leads to a Kalman filter. Using the result in Proposition 3, we obtain a distributed Kalman filter that incorporates the consensus averaging step in (15). To allow correlation between the motion and measurement noise, we follow Crassidis and Junkins [37, Ch. 5] and obtain a correlated Kalman filter in Appendix A-C.
@@ V Distributed MSCKF
In this section, we use Proposition 3 to derive a distributed version of the MSCKF algorithm [8], summarized in Algorithm 1. Each step of the algorithm is described in the following subsections.
@@ V-B Consensus averaging
which is the same as (15) except that we only consider one time step $t$ here. Then, we need to reconstruct the new joint distribution $\mathcal{N}(\bar{\boldsymbol{\mu}}_{i,t-1},\bar{\Sigma}_{i,t-1})$. Since we store the historical camera poses as $SE(3)$ matrices, Proposition 2 can not be applied directly. Following [36, Ch. 7.3.1], we define a Gaussian distribution over a historical camera pose $\underline{T}_{i,k},\ k=t-c,\cdots,t-1$ by adding a perturbation $\epsilon_{i,k}$:
@@
where $(\cdot)^{\wedge}$ defined in [36, Ch. 7.1.2] converts from $\mathbb{R}^{6}$ to a $\mathbb{R}^{4\times 4}$ twist matrix. The estimated covariance $\Sigma_{i,t-1}$ already takes account of both the poses and landmarks. For consistency of notation with Proposition 2, we denote the mean of the pose perturbation as $\boldsymbol{\mu}^{\mathbf{x}}_{i,t-1}=\mathbf{0}_{6c}$. After averaging and reconstructing the new joint distribution, $\bar{\boldsymbol{\mu}}^{\mathbf{x}}_{i,t-1}$ may be non-zero, so we need to correct the camera poses as follows:
@@ V-C State propagation
where $Ad(\cdot)$ is the adjoint of an $SE(3)$ matrix [36, Chapter 7.1.4], $\mathbf{e}_{6n}\in\mathbb{R}^{6n+3m_{t}}$ is a vector with the $6n$-th element as $1$ and the rest as 00, and $W_{i}\in\mathbb{R}^{6\times 6}$ is the odometry measurement covariance.
@@ V-D State update
Finally, by concatenating the residuals and Jacobians of both geometric and object features, we get the overall residual $\mathbf{r}_{i,t}\!=\![\mathbf{r}^{g\top}_{i,t}\ \mathbf{r}^{o\top}_{i,t}]^{\top}$ and Jacobian $H_{i,t}\!=\![H^{g\top}_{i,t}\ H^{o\top}_{i,t}]^{\top}$. As shown in Appendix A-C and [37, Ch. 5], the Kalman gain is:
@@ A-C Derivation of the Kalman filter with noise correlation
Following the approach in Eq. 3.113 in [36, Chapter 3.3.2], we multiply both sides of (50) by the matrix:
@@
which aligns with [37, Table 5.1].